OpenStudy (anonymous):

1. Evaluate the expression. 2(y + 2) for y = 3 (1 point) –4 8 10 2 2. The cost in dollars of a school banquet is 74 + 12n, where n is the number of people attending. What is the cost for 52 people? (1 point) $550 $698 $624 $138 3. Write a word phrase for p ÷ 3. (1 point) a number plus 3 the product of a number and 3 the quotient of a number and 3 a number decreased by 3 4. A major league baseball player chews 17 pieces of gum per game. Write an algebraic expression to show how many pieces of gum he mig

OpenStudy (anonymous):

don't answer the 4th one

OpenStudy (anonymous):

I mess up

OpenStudy (anonymous):

plz help

OpenStudy (anonymous):

1. 2([3]+2) = 6+4 = 10 2. 74 + 12(52) = $624 3. The quotient is the result of division. Therefore, "the quotient of a number and 3" would be the correct answer. Do you understand how I determined the answers to the questions?